OpenStudy (anonymous):

Two numbers have a sum of 9 and a product of 20. What single-variable quadratic equation in the form ax^2 + bx + c = 0 can be used to represent the product of the two numbers?

Parth (parthkohli):

Let's say that x is number 1. The second number would be (9 - x) Now, can you do it? =)

OpenStudy (anonymous):

X + Y = 9 X * Y = 20

Parth (parthkohli):

But it's written -SINGLE VARIABLE-

OpenStudy (anonymous):

What does that mean\?

Parth (parthkohli):

You cannot use two variables.

OpenStudy (anonymous):

Weird

Parth (parthkohli):

The two variables that you can use are x and (9 - x).

OpenStudy (anonymous):

X - 9 uses 1 variable right

Parth (parthkohli):

Yes.

Parth (parthkohli):

We're using only one variable, i.e x.

OpenStudy (anonymous):

X - 9 why does that work?

Parth (parthkohli):

Because if one number is x, and the second number and x add up to 9, then we can use 9 - x, rite?

OpenStudy (anonymous):

What about X - 9?

Parth (parthkohli):

Okay, that wouldn't work. I know that the answer is 4 and 5. Take 4 as x. Would x - 9 work?

OpenStudy (anonymous):

No

Parth (parthkohli):

Exactly. So here's your answer.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

@ParthKohli I do not know where to go from 9 - X

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow x + (9 - x) = 9 }\) \(\Large \color{MidnightBlue}{\Rightarrow x + 9 - x = 9 }\) \(\Large \color{MidnightBlue}{\Rightarrow 0 = 0 }\) Agree?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Then: \(\Large \color{MidnightBlue}{\Rightarrow x * (9 - x) = 20 }\) \(\Large \color{MidnightBlue}{\Rightarrow x * (9 - x) - 20 = 0 }\) \(\Large \color{MidnightBlue}{\Rightarrow -x^2 +9x - 20= 0 }\) And here is your quadratic equation.

OpenStudy (anonymous):

@ParthKohli I lied ParthKohli when I said I agree'd

Parth (parthkohli):

Lol

OpenStudy (anonymous):

9 - x well x could resemble 4 right?

OpenStudy (anonymous):

So 9 - 4 = 5 then 5 + 4 = 9 again right?|

Parth (parthkohli):

Sure.

OpenStudy (anonymous):

Is that the wrong way to think about it?

Parth (parthkohli):

No, I guess

OpenStudy (anonymous):

Okay so let's pretend x is 4 then

OpenStudy (anonymous):

x∗(9−x)=20 well it's like 9 - 4 = 5, 5 * 4 = 20

OpenStudy (anonymous):

Right?